OpenStudy (anonymous):

evaluate lim x over square root x^2 +1, x-> infinity.

OpenStudy (anonymous):

= lim x> inf, sqrt[( x^2 /( x^2+1)] = lim x> inf, x^2 /( x^2+1) using l'hopitals rule = lim x>inf. 2x / 2x = 1